Please help translate this sentence. Mr. Lath rom offered to pay Karen 4 times as much to rake leaves this year than he offered her last year. Karen receives $30 this year. How much did Mr. Lathrom pay last year?
Answer by Maths68(1474)

Let
Last year Karen received = x
This year Karen received = 4x
This year Karen received = $30
4x=30
4x/4=30/4
x=15/2
x=$7.5
Mr. Lothrom paid last year = $7.5